OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = 9 –3x + 2y = 9

OpenStudy (anonymous):

add both the equations it will eliminate x then find y after that put the value of y to any equation to find x

hero (hero):

The goal is to help lead the student who posted the question to the answer. So if you provided these steps, and he didn't get it the first time, the next time he posts a question, you should provide more explanation and more steps.

hero (hero):

The goal is to be willing to help the student until he or she "gets it".

OpenStudy (anonymous):

3x - 4y = 9 9 - 3x + 2y = 9

OpenStudy (anonymous):

3x - 4y =9 + 3x + 2y = 0

OpenStudy (anonymous):

so we get 6y = 9 then y = 3/2

OpenStudy (anonymous):

3x - 4y = 9 3x - 4(3/2) = 9 so x = -1